Eric J. P. Mumpower

nocturne@mit.edu

Term Address 259 St. Paul Street Brookline, MA 02146

(617) 734-9211 Home Address 4133 Thornwall Dr. Marietta, GA 30062 (404) 973-5133

Professional Interests

Topics related to distributed information systems and the World Wide Web.

Experience MIT Department of Computer Science/Electrical Engineeering Cambridge, MA Computation Structures (6.004) Lab Assistant February 1994 - Present Responsible for instructing and assisting students in construction and debugging of their ‘MAYBE’ computers, then testing their understanding of the machine for lab check-off. Worked only during regular (fall and spring) academic terms.

MIT Student Information Processing Board Cambridge, MA Volunteer Webmaster January 1994 - Present Responsible for a portion of the maintainence and administration of the web server www.mit.edu. Implemented a variety of administrative tools and CGI scripts.

ISSC (wholly owned subsidiary of IBM) Lakeside facility, Atlanta, GA Helpdesk Automation Assistant June-August, 1993 Automated HelpDesk procedures for technical support staff. Updated and expanded all software components in a VM/CMS-to-digital-pager gateway used throughout the Lakeside facility.

ISSC Hillside facility, Atlanta, GA Network Automation Assistant June-August, 1992 Assisted in automation of IBM network management services for networks local to Atlanta. Learned several languages for creation and modification of automation software.

Education MASSACHUSETTS INSTITUTE OF TECHNOLOGY Cambridge, MA Completing B.S. in Computer Science in May, 1996

Skills Computer Languages: C, C++, Perl, CLU, Lisp, Scheme, HTML, LaTEX, generic microcode and nanocode, some assembly code

Operating Systems: UNIX, MS-DOS, OS/2

Relevant Coursework  6.001, 6.034, 6.270, 6.170, and 6.004. Currently registered for 6.033 and 6.042.

Background/ Interests SIPB Member, National Merit Finalist UNIX hacking, pottery, music composition.

Awards Computation Structures (6.004) Design Contest (“Largest speedup in the x-y plane”) Was on a 2-person team which made a 6.004 lab kit independently mobile. The kit was equipped with 6.270 motors and sensors, which were controlled by the same circuitry and microcode which ran the competition code, and thus wall-followed its way around the room as it competed in the design contest.